### v1

# Charming Darwin: the evolution of QCD parameters across different species

Abstract
We explore the prospects of measurements of spectral moments of inclusive charm decays with BESIII. The rich and uniquely clean data set of charm mesons and baryons at BESIII offers a unique laboratory to study the evolution of Heavy Quark Expansion (HQE) parameters across different charm hadron species and to shed light on the interplay between heavy quark dynamics and light quark effects. The HQE in terms of inverse powers of the heavy meson mass is well-established in beauty decays, however due to the lighter charm mass its applicability to charm remains an open question. To date no determination of the HQE parameters, the kinetic energy, chromomagnetic moment and Darwin terms, has been attempted. A particular important role here is given to the Darwin and weak annihilation operators, whose values are important to predict lifetimes of heavy hadrons. Using a fast simulation for the BESIII detector response and predictions for spectral moments, we investigate the sensitivity to HQE parameters with today’s and possible future data sets. In addition, we discuss the theory challenges for the HQE in charm and the experimental limitations. We further investigate the sensitivity of determining the CKM matrix element $|V_{cs}|$ with inclusive semileptonic charm decays.
1 Introduction
The Heavy Quark Expansion (HQE) is a powerful technique to study the dynamics of inclusive heavy hadron decays and predict their lifetimes. In this formalism, the decay rates of heavy quarks can be related to the forward scattering amplitude and the time ordered products can be expanded using the Operator Product Expansion (OPE). The first non-trivial contribution to the decay rate arises from dimension-5 operators (at order $1/m_{Q}^{2}$), where $m_{Q}$ is the mass of the heavy quark and the non-perturbative matrix elements $\{\mu_{\pi}^{2},\mu_{G}^{2}\}$ correspond the kinetic energy and the chromomagnetic moment of the heavy quark inside the heavy meson. Dimension-6 operators (entering at order $1/m_{Q}^{3}$) result in two additional contributions $\{\rho_{D}^{3},\rho_{\mathrm{LS}}^{3}\}$ corresponding to the Darwin and spin-orbit terms. The Darwin operator provides one of the key corrections to the free heavy quark decay lifetime Lenz:2020oce and the HQE is constructed in such a way that it depends on the flavour of the initial heavy hadron.
For inclusive semileptonic beauty meson decays, the HQE has achieved remarkable precision, with hadronic matrix elements identified up to dimension 8 Mannel:2023yqf and $\alpha_{s}^{3}$ corrections to the total decay rate calculated Fael:2020tow . This progress, combined with experimental precision, enables the extraction of inclusive $|V_{cb}|$ at the percent level (see, e.g., Bernlochner:2022ucr ; Finauri:2023kte and references therein). Furthermore, the possibility of measuring inclusive $B_{s}$ decays using a sum-over-exclusive technique has been studied at the LHC DeCian:2023ezb .
For charm decays, the HQE is expected to converge more slowly due to the much smaller expansion scale $\Lambda_{\rm QCD}/m_{c}$. Its validity has been debated over the years Voloshin:2002je ; Falk:1995kn ; Lenz:2013aua ; Lenz:2014jha , with a recent reevaluation in Fael:2019umf . Precisely due to the slower convergence, inclusive charm decays are significantly more sensitive to HQE hadronic matrix elements, such as $\rho_{D}^{3}$, as well as to weak annihilation (WA) operators. These WA operators are crucial input parameters for Standard Model predictions of $B\to X_{u}\ell\bar{\nu}_{\ell}$  Ligeti:2010vd ; Becirevic:2008us ; Gambino:2010jz ; Bigi:2009ym and $B\to X_{s,d}\ell\ell$ Huber:2015sra ; Huber:2019iqf ; Huber:2020vup decays, as well as for lifetime predictions Albrecht:2024oyn . Charm meson and baryon lifetimes have recently been revisited and measured with thus far unprecedented precision by Belle II Belle-II:2021cxx ; Belle-II:2022ggx ; Belle-II:2023eii ; Belle-II:2022plj .
In this paper, we explore the opportunities that the first set of measurements of inclusive semileptonic charm decays could provide. We assess the potential of the large data set recorded by the Beijing Electron Spectrometer Mk. III (BESIII), which is particularly intriguing due to its clean samples of $D^{0}$, $D^{+}$, $D_{s}^{+}$, and $\Lambda_{c}^{+}$ decays. Such an experimental program, combined with corresponding theoretical progress, would a permit a detailed study of the evolution of non-perturbative QCD parameters across various species of heavy hadrons, including the first determination of these parameters in heavy baryon decays. In addition, this program would, for the first time, link semileptonic charm hadron measurements with precision tests of the Standard Model through charm hadron lifetimes.
Previous measurements of inclusive semileptonic $D^{0}$, $D^{+}$, and $D_{s}^{+}$ decays at CLEO CLEO:2009uah have already been used to determine the magnitude of weak annihilation operators Ligeti:2010vd ; Gambino:2010jz . More recently, measurements from BESIII BESIIIDsIncSL ; BESIII:2022cmg have been used to determine the strength of the strong coupling constant Wu:2024jyf at the charm scale.
In Ref. Gambino:2010jz , CLEO’s differential measurements of lepton energies were converted into moments. When combined with the HQE parameters obtained from $B\to X_{c}\ell\bar{\nu}_{\ell}$ inclusive decays, this enabled the extraction of weak annihilation effects, which were found to be small, suggesting that the HQE may converge sufficiently rapidly. We propose a dedicated analysis of the lepton energy moments and the as-yet unexplored di-lepton invariant mass moments of $D^{0}$, $D^{+}$, $D_{s}^{+}$, and $\Lambda_{c}^{+}$ decays. This approach could pave the way for a full determination of charm HQE parameters directly from charm data. The comparison between $D$ and $D_{s}^{+}$ decays would allow for an $SU(3)$ symmetry test based on data, providing crucial inputs for lifetime determinations Lenz:2020oce . Moreover, comparing these findings with the HQE parameters obtained from inclusive $B$ decays will be intriguing. Finally, combining these results with branching ratio measurements CLEO:2009uah ; BESIIIDsIncSL ; BESIII:2022cmg could facilitate the extraction of both $|V_{cs}|$ and $|V_{cd}|$ in a manner analogous to the inclusive $|V_{cb}|$ extraction Bernlochner:2022ucr ; Bordone:2021oof ; Finauri:2023kte , with the potential to achieve uncertainties competitive with exclusive determinations.
The remainder of this manuscript is organized as follows: Section 2 introduces a realistic fast simulation of BESIII for semileptonic decays, including estimates for systematic uncertainties. Section 3 discusses the experimental approach to calibrating detector-level observables to unfolded raw spectral moments. Section 4 reviews the HQE for charm and details the challenges associated with the lighter charm quark compared to beauty decays. Section 5 presents an exploratory study on extracting the experimental precision of HQE parameters from the unfolded spectral moments. Lastly, Section 6 provides a detailed outlook on the prospects for the inclusive charm program outlined in this work.
2 Prospects for inclusive semileptonic measurements at BESIII
BESIII records electron-positron collisions provided by the Beijing Electron-Positron Collider Mk. II (BEPCII). BEPCII can deliver center-of-mass energies between 1.85 and 4.95 GeV and the BESIII experiment’s large data sets at open-charm-hadron pair-production thresholds provide ideal data samples to perform inclusive semileptonic measurements in the charm system. From the interior to the exterior, the BESIII detector BESIIIDetector is comprised of a drift chamber tracking system, a plastic scintillator time-of-flight system, a crystal calorimeter, a 1T super-conductor solenoid, and a resistive-plate-chamber muon system. The drift chamber provides sub-percent level momentum resolution of charged particles and, combined with measurements from the time-of-flight system, is able to distinguish well between protons, charged pions, and charged kaons. The BESIII calorimeter allows for excellent identification of electrons and provides percent-level resolution on energy depositions of charged and neutral particles.
The sensitivities to HQE parameters from measurements of $D^{+}\to Xe^{+}\nu_{e}$, $D^{0}\to Xe^{+}\nu_{e}$, $D_{s}^{+}\to Xe^{+}\nu_{e}$, and $\Lambda_{c}^{+}\to Xe^{+}\nu_{e}$ at BESIII are estimated by analyzing Monte Carlo simulation produced with the EvtGen package EvtGen . A measurement strategy similar to that employed in inclusive semileptonic beauty decays at Belle and Belle II BelleQ2 ; BelleIIQ2 is assumed, in which the lepton and all hadrons accompanying the lepton in the final state of the semileptonic decay are reconstructed. This differs from techniques applied in previous measurements of inclusive semileptonic charm decays CLEO:2009uah ; BESIIIDsIncSL ; BESIII:2022cmg , in which the hadronic system $X$ was not reconstructed and only the charged lepton was identified. These measurements produced precise determinations of the inclusive branching fractions and the electron momentum spectra in the laboratory frame. With additional assumptions on the production mechanism of charmed-hadron pairs, analysis of the measured laboratory-frame distributions allowed for HQE analysis of the $D^{0}$, $D^{+}$, and $D_{s}^{+}$ decays to constrain the contribution of weak-annihilation operators Gambino:2010jz ; Ligeti:2010vd . In this study, we examine how precisely the explicit reconstruction of the hadronic $X$ system can be performed to construct Lorentz-invariant observables. Such quantities are crucial for mitigating the proliferation of HQE parameters at higher orders Fael:2018vsp ; Mannel:2018mqv . This reconstruction enables the reconstruction of the Lorentz-invariant lepton-neutrino mass squared, $q^{2}$, and the Lorentz-invariant mass of the hadronic system, $M_{X}$. Additionally, properties of the $X$ system allow for the separation of $c\to s$ and $c\to d$ transitions, facilitating the first extraction of the CKM matrix elements $|V_{cs}|$ and $|V_{cd}|$ from inclusive semileptonic charm decays.
In the explored analysis procedure, double-tag techniques similar to those developed by the MARKIII collaboration MARKIII_one ; MARKIII_Two are employed with the tag decay modes [^1] Charge conjugation is implied here and throughout the paper. $\overline{D}{}^{0}\to K^{-}\pi^{+}$, $D^{-}\to K^{-}\pi^{+}\pi^{+}$, from Ref. BESIII:2023cym for $D_{s}^{-}$, and from Ref. BESIII:2022cmg for $\Lambda_{c}^{-}$. The dataset listed for BESIII analysis of $D_{s}^{+}$ mesons primarily contains $D_{s}^{+}$ mesons produced through $e^{+}e^{-}\to D_{s}^{*+}D_{s}^{-}$, and so the reconstruction of a $D_{s}^{*}$ meson through a $D_{s}^{*}\to\gamma D_{s}$ is also assumed. The assumed size of BESIII data samples are listed Table 1. The listed data sets correspond to samples which BESIII has analyzed for publication, with the exception of the samples for $D^{0}$ and $D^{+}$ analysis, where a larger dataset has recently been collected.
Table 1: Integrated luminosities, center-of-mass energies and estimated inclusive semileptonic double-tag yields of BESIII data sets assumed in sensitivity studies.
$D^{0}$ $D^{+}$ $D_{s}^{+}$ Ds_Lumi $\Lambda_{c}^{+}$ Lc_Lumi1 ; Lc_Lumi2
[…]
Estimated Double-tag Yields 200000 700000 30000 4300
2.1 Estimated distributions from fast simulation
A fast-simulation software based on EvtGen EvtGen is developed to estimate the effects of BESIII detector response, momentum resolution, and geometric and kinematic acceptance on inclusive semileptonic measurements. The effects of final-state radiation are neglected in the simulation. The decays of charmed hadrons are generated by EvtGen for the tag-side meson and all observed semileptonic decays PDG2022 , and those predicted by isotopic symmetry based on observed decays. The ISGW2 ISGW2 model is assumed for all $\{D^{+},D^{0},D_{s}^{+},\Lambda_{c}^{+}\}\to Me^{+}\nu$ decays, where $M$ is any hadron, and a uniform phase-space model is assumed when more than one hadron is produced in the semileptonic decay. Detector effects are incorporated probabilistically based on publicly reported reconstruction efficiencies and resolutions. The efficiencies of tracking charged particles and particle identification are estimated based on the published measurement of the branching fraction of $D_{s}^{+}\to Xe^{+}\nu_{e}$ decays BESIIIDsIncSL . Photon reconstruction efficiencies and resolution are estimated based on Ref. BESIIINeutrals . The energy deposition of $K^{0}_{L}$ mesons and neutrons in the BESIII calorimeter is estimated based on the analysis of $K^{0}_{L}$ backgrounds in the search for $D^{0}\to\pi^{0}\nu\overline{\nu}$ at BESIII BESIIIPiNuNu . The reconstruction efficiency and momentum resolution of $K_{S}^{0}\to\pi^{+}\pi^{-}$ decays is estimated based on Ref. BESIIIIncLKS . The efficiency and momentum resolution of reconstructing $\Lambda\to p\pi^{-}$ decays is assumed to be identical to $K_{S}^{0}\to\pi^{+}\pi^{-}$ decays due to the similar vertex reconstruction.
Simulated measured distributions of the kinematic distributions of interest in the HQE (the squared four-momentum of the $e^{+}\nu$ system $q_{\text{reco}}^{2}$, the $e^{+}$ energy in the parent rest frame $E_{\ell}$, and the mass of the hadronic system in the final state $M_{X}$) are produced by the fast simulation based on the estimated double-tag sample sizes listed in Table 1. The BESIII detector is unable to consistently distinguish between electrons and other charged particles for candidate tracks with momenta less than 200 MeV. As such, it is required that the momentum of the electron candidate be above this threshold. The resulting $E_{\text{miss}}-p_{\text{miss}}$ distribution produced by the fast simulation of for each charmed hadron is shown in Fig. 1, with the contribution of decays with $K^{0}_{L}$ mesons or $K^{0}_{S}\to\pi^{0}\pi^{0}$ decays in the final state highlighted. As the figure shows, decays with $K^{0}_{L}$ mesons or $K^{0}_{S}\to\pi^{0}\pi^{0}$ are often poorly reconstructed, due to the poor resolution of $K^{0}_{L}$ energy depositions in the calorimeter and the inability of the BESIII detector to distinguish between energy depositions due to photons and those due to $K^{0}_{L}$ mesons. As such, a requirement of $\left|E_{\text{miss}}-p_{\text{miss}}\right|<500\text{ MeV}$ is placed on simulated events to estimate the sensitivity of measurements at BESIII, and a correction is later applied to account for this requirement. It should be noted that a measurement at BESIII could correct for decays with $K^{0}_{L}$ mesons or $K^{0}_{S}\to\pi^{0}\pi^{0}$ decays in the final state that are removed with this selection based on measurements of decays with $K^{0}_{S}\to\pi^{+}\pi^{-}$ decays in the final state, which BESIII reconstructs with high efficiency and good momentum resolution. Contributions from neutrons, primarily in $\Lambda_{c}^{+}$ decays, can similarly be corrected for assuming isotopic symmetry [^2] The decay $\Lambda_{c}^{+}\to ne^{+}\nu$ must be handled separately. based on observed decays with protons in the final state.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 1: Distribution of the reconstructed $E_{\text{miss}}-p_{\text{miss}}$ for decays of (top-left) $D^{0}\to Xe^{+}\nu_{e}$, (top-right) $D^{+}\to Xe^{+}\nu_{e}$, (bottom-left) $D_{s}^{+}\to Xe^{+}\nu_{e}$, and (bottom-right) $\Lambda_{c}^{+}\to Xe^{+}\nu_{e}$ produced by the fast simulation. The light blue histogram is the total distribution, and the dark blue histogram corresponds to events with $K^{0}_{L}$ mesons, $K^{0}_{S}\to\pi^{0}\pi^{0}$ decays, or neutrons in the final state. The $E_{\text{miss}}-p_{\text{miss}}$ selection requirements correspond to the interior of the dashed red lines.
The kinematic and geometric acceptance of BESIII associated with reconstructing and identifying the positron and the $E_{\text{miss}}-p_{\text{miss}}$ requirement reduce the sample size for analysis, and sculpt the kinematic distributions of interest. The efficiency of by the acceptance and selections is approximately 68%, 53%, 44%, and 49% for the $D^{0}\to Xe^{+}\nu_{e}$, $D^{+}\to Xe^{+}\nu_{e}$, $D_{s}^{+}\to Xe^{+}\nu_{e}$, and $\Lambda_{c}^{+}\to Xe^{+}\nu_{e}$ samples, respectively. While the resolution is predicted to be considerably worse for the $D_{s}^{+}$ sample, a measurement at BESIII could improve this with a kinematic constraint on the $D_{s}^{*}$ candidate. Comparisons between the generated kinematic distributions of interest of $D^{+}\to Xe^{+}\nu_{e}$ decays and the corresponding reconstructed distributions produced by the fast simulation are shown in Fig. 2.
Refer to caption Refer to caption
Figure 2: Comparisons between the generated $D^{+}$ distributions of the (top-left) $E_{\ell}$, (top-right) $M_{X}$, and (bottom) $q_{\text{reco}}^{2}$ variables and the reconstructed distributions produced by the fast simulation. In each plot, the generated distribution is shown as the dashed orange histogram and the reconstructed distribution is shown as the solid light blue histogram.
After the selection requirements are applied, the event-by-event difference between reconstruction and generated variables exhibit a clear core resolution with a tail due to missing particles in the reconstructed $X$ system. The effects of the BESIII resolution can be corrected for with a calibration procedure similar to Refs. BelleQ2 ; BelleIIQ2 . The distribution of the difference of the reconstructed and generated $q^{2}$, referred to as $\Delta q^{2}$, is shown in Fig. 3, where all discussed selections have been applied.
$D^{0}\to Xe^{+}\nu_{e}$ $D^{+}\to Xe^{+}\nu_{e}$
Refer to caption Refer to caption
$D_{s}^{+}\to Xe^{+}\nu_{e}$ $\Lambda_{c}^{+}\to Xe^{+}\nu_{e}$
Refer to caption Refer to caption
Figure 3: Predicted distribution of $\Delta q^{2}$ for events passing selections from (top-left) $D^{0}\to Xe^{+}\nu_{e}$, (top-right) $D^{+}\to Xe^{+}\nu_{e}$, (bottom-left) $D_{s}^{+}\to Xe^{+}\nu_{e}$, and (bottom-right) $\Lambda_{c}^{+}\to Xe^{+}\nu_{e}$.
The differentiation of inclusive decays proceeding through $c\to s$ and $c\to d$ transitions can be achieved by identifying the total strangeness of the hadronic system $X$. This can be determined from the number of observed $K^{+}$ mesons, $K_{S}^{0}\to\pi^{+}\pi^{-}$ decays, and $\Lambda\to p\pi^{-}$ decays with high efficiency. It should be noted again that the requirements on $E_{\text{miss}}-p_{\text{miss}}$ suppress contributions from decays with $K_{L}^{0}$ mesons, other $K_{S}^{0}$ decays, or neutrons in the final state, but that this can be corrected based on the observed decays with reconstructed $K_{S}^{0}\to\pi^{+}\pi^{-}$ and proton candidates. Estimates of the capability of $c\to s$-tagging is shown in Fig. 4. The residual contribution of events where the strangeness of the final state is misidentified (responsible for the large contribution of misidentified events in the $c\to d$ distribution) could be determined through a simultaneous template fit of the $M_{X}$ distributions for the $c\to s$-tagged sample and the $c\to d$-tagged sample.
Refer to caption Refer to caption
Figure 4: Estimated $M_{X}$ distribution of (left) $c\to s$-tagged $D^{+}$ semileptonic decays and (right) $c\to d$-tagged $D^{+}$ semileptonic decays. In each case, the expected observed distribution is shown in light blue, the contribution of events with $K^{0}_{L}$ mesons and $K_{S}^{0}\to\pi^{0}\pi^{0}$ decays in the final state is shown in dark blue, and the true distribution of events passing selections shown as the dashed orange histogram.
2.2 Systematic uncertainties
The size of systematic uncertainties in BESIII measurements of the differential spectra due to detector modelling are estimated by varying the assumed reconstruction resolution and the reconstruction and particle identification efficiencies, within their stated uncertainties and repeating the procedure described in Sec. 2.1. The calibration procedure described in Sec. 3 depends on the branching fractions that are assumed for the component exclusive branching fractions. Alternate spectra are produced based on the measured uncertainties of the exclusive branching fractions included in the fast simulation and checked to determine the magnitude of the dependence.
2.3 Background contributions
While we consider all foreseen detector effects on reconstructing and identifying signal events in the fast simulation, we do not consider possible contributions from backgrounds. Backgrounds will arise from the tag-side hadron being misidentified, the lepton candidate being a misidentified hadron, and contributions from $e^{+}e^{-}\gamma$ Dalitz decays of light mesons produced in the final state of charmed hadron decays. The first category can be handled in a similar fashion to the inclusive semileptonic beauty measurements at Belle and Belle II BelleQ2 ; BelleIIQ2 . The second category can be handled through a weighting procedure, where particle identification observables are used to weight the observed sample. The third and final category can be corrected for with a similar procedure to previous inclusive semileptonic charm measurements CLEO:2009uah ; BESIIIDsIncSL ; BESIII:2022cmg , where these backgrounds are corrected for with data control samples in which the identified lepton has opposite charge to the expected flavour of the parent hadron. The expected impact of background subtraction on the statistical precision are estimated from the purities observed in the previous inclusive semileptonic charm measurements CLEO:2009uah ; BESIIIDsIncSL ; BESIII:2022cmg , described in more detail in Sec. 5.
3 Calibration procedure and sensitivity of raw spectral moments
We use the simulated events to determine the raw moments of $q^{2}$ and $E_{\ell}$ of order $n$ and for a progression of threshold cuts on $q^{2}$. This is done in a three-step procedure following the approach from Ref. BelleQ2 ; BelleIIQ2 . First, we apply a linear calibration as a function of $q^{2n}$ or $E_{\ell}^{n}$, exploiting the fact that the reconstructed spectral moments of the distributions of interest are mostly linearly shifted with respect to their true values. We determine the slope $c_{n}$ and intercept $m_{n}$ of the linear calibration function via a fit to the simulated samples that solves
$q^{2n}_{\text{cal}}$ $=\left(q^{2n}_{\text{reco}}-c_{n}\right)/m_{n}\,,$ $E_{\ell,\text{cal}}^{n}=\left(E_{\ell,\text{reco}}^{n}-c_{n}\right)/m_{n}\,.$ (1)
Here the subscript ‘reco’ indicates the reconstructed moments and ‘cal’ indicates the calibrated moments. We further derive individual calibration constants for $q^{2n}$ and $E_{\ell}^{n}$ and for each order $n$. In Figure 5 we show the linear calibration function for the first $q^{2}$ moment for the $D^{0}$ decay: the reconstruted and generated values are approximately linearly shifted.
Figure 5: The linear calibration function determined for the first $q^{2}$ moment of the $D^{0}$ decay which is used to calibration the reconstructed moments to the corresponding generator moments.
Second, we correct the calibrated moments for deviations from the linear calibration function via a correction factor derived from simulation,
$C_{\mathrm{calib}}$ $=\langle q^{2n}_{\mathrm{gen,sel}}\rangle/\langle q^{2n}_{\mathrm{cal}}\rangle\qquad\text{or}\qquad\langle E^{n}_{\ell,\mathrm{gen,sel}}\rangle/\langle E^{n}_{\ell,\mathrm{cal}}\rangle\qquad,$ (2)
where the subscript ‘gen,sel’ indicates the generator level moments with selections applied. Third, we correct for the selection and acceptance effect with a correction factor, also derived from simulation,
$C_{\mathrm{gen}}$ $=\langle q^{2n}_{\mathrm{gen}}\rangle/\langle q^{2n}_{\mathrm{gen,sel}}\rangle\qquad\text{or}\qquad\langle E^{n}_{\ell,\mathrm{gen}}\rangle/\langle E^{2n}_{\ell,\mathrm{gen,sel}}\rangle\qquad$ (3)
where the subscript ‘gen’ indicates the generator level moments without any selections or acceptance effects imposed. Both $\mathcal{C}_{\mathrm{calib}}$ and $\mathcal{C}_{\mathrm{gen}}$ are determined for each $q^{2}$ threshold and shown in Figure 6. The $\mathcal{C}_{\mathrm{calib}}$ factors range between 0.97 and 1.03 for the $\langle q^{2n}\rangle$ and are very close to unity for the $\langle E_{\ell}^{n}\rangle$. The $\mathcal{C}_{\mathrm{gen}}$ factors range between 0.75 and 1.00 for both $\langle q^{2n}\rangle$ and $\langle E_{\ell}^{n}\rangle$. The origin of these large factors are decays that are only partially reconstructed due to the presence of e.g. $K_{L}$, and are rejected by the $|E_{\mathrm{miss}}-p_{\mathrm{miss}}|>0.5\,\mathrm{GeV}$ selection (cf. Fig. 1). In Figure 6 we show the calibration factors for the first $q^{2}$ moment for the $D^{0}$ decay at different thresholds $q^{2}_{\mathrm{th}}$.
Figure 6: The $\mathcal{C}_{\mathrm{calib}}$ (left) and $\mathcal{C}_{\mathrm{gen}}$ (right) calibration factors for the $D^{0}$ decay. The former corrects for deviations from the linear calibration function, and the latter corrects for selection and acceptance effects.
Both $\mathcal{C}_{\mathrm{calib}}$ and $\mathcal{C}_{\mathrm{gen}}$ exhibit a similar behaviour among the four decays considered. A more detailed discussion on the linear calibration functions and the multiplicative calibration factors can be found in Appendix A.
Applying all outlined calibration steps, we determine the spectral moments via
$\langle q^{2n}\rangle$ $=\sum_{i}q^{2n}_{i\,\text{cal}}\times C_{\mathrm{calib}}\times C_{\mathrm{gen}}\,,$ $\langle E_{\ell}^{n}\rangle$ $=\sum_{i}E^{n}_{\ell,i,\text{cal}}\times C_{\mathrm{calib}}\times C_{\mathrm{gen}}\,,$ (4)
with $i$ denoting a given event. The absolute precision, including both statistical and systematic uncertainties, that can be achieved for the moments based on our simulation outlined above is listed in Table 2 and shown in Figure 7.
Table 2: Relative uncertainties in % on the determined moments with the calibration procedure described in the text. The uncertainties include statistical and systematic uncertainties. We only show the uncertainty for the moments with $q^{2}_{\mathrm{th}}=0$.
$\langle q^{2}\rangle$ $\langle q^{4}\rangle$ $\langle q^{6}\rangle$ $\langle q^{8}\rangle$ $\langle E_{\ell}^{1}\rangle$ $\langle E_{\ell}^{2}\rangle$ $\langle E_{\ell}^{3}\rangle$ $\langle E_{\ell}^{4}\rangle$
$D^{0}$ $1.2$ $3.0$ $6.1$ $11.3$ $0.3$ $0.4$ $0.5$ $0.6$
$D^{+}$ $0.6$ $0.8$ $1.2$ $1.6$ $0.2$ $0.3$ $0.4$ $0.6$
$D_{s}$ $10.5$ $11.7$ $9.8$ $8.1$ $0.9$ $1.7$ $2.9$ $10.4$
$\Lambda_{c}$ $5.6$ $9.9$ $15.0$ $18.7$ $4.2$ $5.9$ $7.2$ $8.3$
Figure 7: Relative uncertainties in % on the determined moments with the calibration procedure described in the text. The uncertainties include statistical and systematic uncertainties. We only show the uncertainty for the moments with $q^{2}_{\mathrm{th}}=0$. We also show relative uncertainty on the moments determined by CLEO from Ref. CLEO:2009uah ; Gambino:2010jz .
4 The Heavy Quark Expansion for charm
4.1 Theoretical framework
The setup of the Operator Product Expansion (OPE) for inclusive $D\to X\ell\bar{\nu}_{\ell}$ decays differs from the standard, well-established setup for $B\to X_{c}\ell\bar{\nu}_{\ell}$ decays due to the hierarchy between the charm-quark and light-quark decay products. In the latter case, the ratio $m_{c}/m_{b}$ is fixed when setting up the OPE. However, for $D\to X_{s}$ and $D\to X_{d}$ decays, the $s$ and $d$ quarks need to be integrated out at their respective scales. This introduces four-quark operators (weak annihilation) (see e.g. Bigi:2009ym ; Breidenbach:2008ua for a detailed discussion). As such, the OPE used to describe charm decays resembles more that of the $B\to X_{u}\ell\bar{\nu}_{\ell}$ case. The setup of the HQE is such that the hadronic parameters depend on the mass of the initial state meson. Therefore, the hadronic matrix element differ between charm and beauty decays, but also between $D_{s}$ and $D$ decays.
From the theoretical side, extracting the HQE element from data requires the OPE expressions for the spectral moments at certain order in the $\Lambda_{\rm QCD}/m_{c}$ and $\alpha_{s}(m_{c})$ expansions. Here we follow the approach of Ref. Fael:2019umf , which set up the HQE as an expansion in both $\Lambda_{\rm QCD}/m_{c}$, $\alpha_{s}(m_{c})$ and $m_{s}/m_{c}$ for the $c\to s$ transition. The $c\to d$ expressions can be obtained by taking $m_{s}\to 0$.
In the following, we only work up to order $1/m_{c}^{3}$, although the $1/m_{c}^{4}$ terms are also known Fael:2019umf and could be included in a future analysis. We define the hadronic matrix elements $X$ for the two-quark operators following Fael:2019umf ; Fael:2018vsp ; Mannel:2018mqv
$2m_{D}X$ $\equiv\langle D|O^{2q}_{X}|D\rangle\ ,$ (5)
where equivalent expressions for $D_{s}$ or $\Lambda_{c}$ can be written. The matrix elements are then [^3] These are defined in the Reparametrization invariant (RPI) basis. See Fael:2018vsp for a conversion to the spatial derivative basis.
$2m_{D}\mu_{\pi}^{2}$ $=\langle\bar{c}_{v}(iD^{2})c_{v}\rangle$ (6)
$2m_{D}\mu_{G}^{2}$ $=\langle\bar{c}_{v}(iD_{\alpha})(iD_{\beta})(-i\sigma^{\alpha\beta})c_{v}\rangle$ (7)
and at order $1/m_{c}^{3}$
$2m_{D}\rho_{LS}^{3}$ $=\frac{1}{2}\langle\bar{c}_{v}\left\{(iD_{\alpha}),\left[ivD,iD_{\beta}\right]\right\}(-i\sigma^{\alpha\beta})c_{v}\rangle$ (8)
$2m_{D}\rho_{D}^{3}$ $=\frac{1}{2}\langle\bar{c}_{v}\left[(iD_{\mu}),\left[ivD,iD^{\mu}\right]\right]c_{v}\rangle$ (9)
where we have omitted the initial state $D$ mesons for simplicity but note that the matrix elements depend on the initial state meson. In addition, unlike in the setup for $B\to X_{c}$ inclusive decays, the four-quark (weak annihilation) operators remain in the OPE. Following Ref. Fael:2019umf , we define
$22m_{D}T_{i}(\mu_{\rm WA})\equiv\langle D|O^{4q}_{i}|D\rangle,\quad\mbox{with }i=1,$ 2 . (10)
where $\mu_{\rm WA}$ is the weak annihilation scale. At order $1/m_{c}^{3}$ only two operators contribute
[…]
$O_{2}$ $=(\bar{c}_{v}\gamma^{\mu}P_{L}s)\,(\bar{s}\gamma_{\mu}P_{L}c_{v}).$ (11)
The weak annihilation operators $T_{1}$ and $T_{2}$ mix under renormalization with the Darwin term Breidenbach:2008ua ; Bigi:2009ym . This introduces only one new operator Fael:2019umf
$\tau_{0}=128\pi^{2}\left(T_{1}-T_{2}\right)+8\log\left(\frac{\mu_{\rm WA}^{2}}{m_{c}^{2}}\right)\rho_{D}^{3}.$ (12)
We note that our setup, which follows Fael:2019umf , differs from Gambino:2010jz ; Ligeti:2010vd ; Wu:2024jyf due to the three expansion parameters that we employ and through the definition of the four-quark operators given above. In the latter, the WA matrix elements are named $B_{\rm WA}$. Specifically, our expressions differ in the non-logarithmic terms from those of $b\to c$ due to a difference in the matching of the two-quark operators for the $c\to s$ transition. This subtle point was discussed in Fael:2019umf and overlooked in Gambino:2010jz ; Ligeti:2010vd ; Wu:2024jyf . In addition, converting $\tau_{0}$ to $B_{\rm WA}$ also requires setting the scale $\mu_{\rm WA}=0.8$ GeV.
The operators $T_{1}$ and $T_{2}$ discussed above are non-valence Weak Annihilation (WA) operators, which do not depend on the flavour of the spectator quark. For $D^{0}$ decays, only non-valence WA contributes, however for $D_{s}$ and $D^{+}$ decays the valence WA also enters weighted by the relevant CKM factors (see expressions in Sec. 6 of Fael:2019umf ). These operators are defined through (4.1) by replacing $s\to q$, where $q$ is the valence spectator quark. The non-valence WA contribution is similar for $D^{0}$ and $D^{+}$ but differs for $D_{s}$ decays due to $SU(3)$ breaking effects that also affect the other HQE elements. The valence and non-valence WA could be disentangled by measuring the WA parameters, $\tau_{0}$, from $D^{0},D^{+}$ and $D_{s}$ decays Bigi:2009ym . It is important to extract these WA from inclusive charm decays as they form a crucial input into the $D^{0},D^{+}$ and $D_{s}$ lifetime calculations.
The WA operators have been obtained from the total rates in Ligeti:2010vd . In Gambino:2010jz , the CLEO inclusive charm data were converted to lepton energy moments and used to extract both the non-valence and valence WA operators. As discussed above, their setup differs from Fael:2019umf employed here, making it more challenging to disentangle the effect of the logarithmic $\rho_{D}^{3}$ terms and the WA effects. Moreover, $q^{2}$ moments are currently not available. As such, redoing the extraction of WA combined with the first extraction of the other HQE parameters in charm would be beneficial. There is also quite some progress in determining the HQE elements directly from lattice QCD calculations Gambino:2017vkx ; Gambino:2022dvu ; Nefediev:2024mjk ; Kellermann:2023yec . In the future, it will be intriguing to compare extractions from data with the lattice calculations.
In order to extract the HQE parameters, we define the moments of the spectrum as
$\mathcal{M}_{n}\equiv\frac{1}{\Gamma_{0}}\int(M)^{n}\frac{\text{d}\Gamma}{\text{d}M}\,\text{d}M\ .$ (13)
where $M$ is the lepton energy $E_{\ell}$ or the dilepton invariant mass $q^{2}$ and $\Gamma_{0}$ is the tree-level normalization factor. In the current work, we only consider lepton energy and $q^{2}$ moments. The latter have the advantage that they are reparametrization invariant (RPI) observables that depend on a reduced set of observables. Specifically, $\rho^{3}_{LS}$ and $\mu_{\pi}^{2}$ do not enter the $q^{2}$ moments Mannel:2018mqv ; Fael:2019umf . The integration limits are over the allowed phase space depending on the employed kinematic thresholds on the lepton energy or $q^{2}$. Here we work with normalized moments defined as
$\langle M^{n}\rangle\equiv\frac{\mathcal{M}_{n}}{\mathcal{M}_{0}}\ .$ (14)
In the $B\to X_{c}\ell\bar{\nu}_{\ell}$ analyses, usually centralized moments are defined. For our exploratory study, we fit the raw moments as defined in (14), which can readily be converted to centralized moments. The expressions for the lepton-energy and $q^{2}$ moments at tree level are given in Fael:2019umf . For the $\Lambda_{c}$ decays, the expressions are similar to those for the $D$ decays, except that all $\sigma\cdot G$ terms cancel due to the spin structure and thus $\mu_{G}^{2}$ and $\rho_{LS}^{3}$ do not contribute [^4] Technically, we first have to split off the spatial and time derivatives in $\mu_{G}^{2}$ by converting from the RPI basis as $\mu_{G}^{2}\to(\mu_{G}^{2})^{\perp}-\rho_{D}^{3}/m_{c}-\rho_{LS}^{3}/m_{c}$. We can then set $(\mu_{G}^{2})^{\perp}$ and $\rho_{LS}^{3}$ to zero.. In the following, we consider moments up to $n=4$. We note that especially these higher order moments are very sensitive to the HQE parameters and would suffer from larger theoretical uncertainties due to neglected $1/m_{c}^{4}$ HQE parameters.
Finally, using information on the total decay semileptonic width allows for an extraction of $|V_{cs(d)}|$. At leading order in $\alpha_{s}$, we have
$\frac{\Gamma(D\to X_{s}\ell\bar{\nu}_{\ell})}{\Gamma_{0}}$ $=\left(1-8\rho-10\rho^{2}\right)\left(1-\frac{\mu_{\pi}^{2}-\mu_{G}^{2}}{2m_{c}^{2}}\right)+\left(-2-8\rho\right)\frac{\mu_{G}^{2}}{m_{c}^{2}}+6\frac{\rho_{D}^{3}}{m_{c}^{3}}+\frac{\tau_{0}}{m_{c}^{3}}\,,$ (15)
where $\rho=(m_{s}/m_{c})^{2}$ and $\Gamma_{0}=G_{F}^{2}m_{c}^{5}|V_{cs}|^{2}/(192\pi^{3})$ and equivalent for $D\to X_{d}$ decays by putting $\rho\to 0$ and replacing $|V_{cs}|\to|V_{cd}|$.
4.2 Theory challenges
For a precision determination of the HQE parameters in charm, the setup of the HQE for charm at leading order in $\alpha_{s}$ does not suffice due to the large value of the strong coupling constant at the relevant scale $m_{c}$. Practically this means higher-order $\alpha_{s}$ corrections should be included, while getting a reliable estimate for the missing higher-orders in $\alpha_{s}$ is challenging. Clearly, a simple scale variation (often done for $B$ decays) would render much larger uncertainties. For charm decays, $\alpha_{s}$ corrections for $q^{2}$ and lepton energy moments can be obtained from Fael:2022frj which uses an expansion in $\delta=1-\sqrt{\rho}$ with $\rho\equiv(m_{q}/m_{Q})^{2}$ for the $m_{Q}\to m_{q}$ transition. For massless quarks, the $\alpha_{s}$ corrections can be obtained from Chen:2022wit for $q^{2}$ moments. For the $\alpha_{s}^{2}$ BLM corrections are known for lepton energy moments Gambino:2010jz ; Aquila:2005hq (with lepton energy threshold cuts). For the total rate, $\alpha_{s}^{2}$ corrections are available Melnikov:2008qs . To our knowledge, for $q^{2}$ moments, no dedicated calculations for $c\to s$ transitions are available. They could be obtained by adapting existing calculations for inclusive $b\to c$ semileptonic transitions. For the latter, the full $\alpha_{s}^{2}$ corrections to the $q^{2}$ moments were recently obtained Fael:2024gyw  [^5] Slightly before that the BLM corrections for the $q^{2}$ moments were calculated Finauri:2023kte . For the lepton energy moments, non-BLM corrections are only known for $B$ decays for fixed $m_{c}/m_{b}$ masses and for several kinematical cuts Biswas:2009rb . In addition, for $b\to c$ decays even the $\alpha_{s}^{3}$ corrections to the total rate and kinematic moments without any kinematic cut are available Fael:2020tow ; Fael:2022frj .
In addition to the NNLO (or higher) corrections, a short-distance mass scheme for the charm quark is required to avoid the bad perturbative behavior of the pole mass. In $B\to X_{c}$ decays it is customary to use the kinetic mass Bigi:1994ga ; Bigi:1996si , which relies on a cut of $\Lambda_{\rm QCD}<\mu<m_{Q}$. For charm decays, this makes the allowed window for $\mu$ rather small. Refs. Gambino:2010jz ; Wu:2024jyf use the kinetic scheme for charm with $\mu=0.5$ GeV, however, as discussed in Fael:2018vsp (see also Ligeti:2010vd ), it remains to be seen if $\mu\simeq 0.5$ GeV is sufficiently away from $\Lambda_{\rm QCD}$. Alternatively, one could employ the $\overline{\rm MS}$ scheme at $\mu=m_{Q}$ Czarnecki:1994pu , although it has been argued that this scheme is less suitable for decays of the $b$ and $c$ quarks given that the typical energy release is much lower than $m_{Q}$ Melnikov:2000qh . In Ref. Ligeti:2010vd , the perturbative series of the $1S$ Hoang:1998ng ; Hoang:2005zw and PS scheme Beneke:1998rk , finding a reasonably well behaved series for the former.
In order to set up a full charm semileptonic program, a suitable charm mass scheme should be investigated further. This also applies to the HQE parameters themselves. In the $B\to X_{c}\ell\bar{\nu}_{\ell}$ analysis, the kinetic scheme for the HQE parameters is also employed. This kinetic scheme change introduces additional perturbative corrections that multiply the coefficients of $\mu_{\pi}^{2}$ and $\rho_{D}^{3}$. The effect on the four-quark parameters $\tau_{0}$ remains to be investigated. In addition, this kinetic scheme change of the HQE parameters suffers from the same issue as the kinetic charm mass definition.
As such, the dedicated experimental program in inclusive charm proposed here should be combined with detailed theoretical study of the perturbative corrections to inclusive charm decays.
5 Towards an extraction of the Darwin operator and $|V_{cx}|$
To check the future potential of a inclusive charm program at BESIII, we perform an exploratory study and analyze the $q^{2}$ and $E_{\ell}$ moments. Given the theoretical challenges described above, we consider only the achievable experimental uncertainty on the extracted HQE parameters. Our sample data contains $\{D^{0},D^{+},D_{s},\Lambda_{c}\}\to(X_{s}+X_{d})\ell\bar{\nu}$ decays. As discussed, in the future these samples could be separated as is done for semileptonic $B\to X_{c}$ versus $B\to X_{u}$ decays. For this first study, we simply neglect the $c\to d$ contribution as it is suppressed by $|V_{cd}/V_{cs}|^{2}=0.04$. In addition, we correct for the lepton energy acceptance when applying the $\mathcal{C}_{\mathrm{gen}}$ factor, so that the measured moments have only the quoted $q^{2}_{th}$ threshold selections applied.
Specifically, we work up to order $1/m_{c}^{3}$ in the HQE and only include $\alpha_{s}$ corrections. It is customary to include an additional theoretical uncertainty due to not included higher orders in the HQE (and $\alpha_{s}$ expansion) (see e.g. the discussion in Finauri:2023kte ). We do not include any theory uncertainties in the fit and leave this for future studies. We also do not consider any external inputs for the $\mu_{G}^{2}$ parameter which can in principle be linked to the mass difference of the initial state and its first excited state.
For the strange quark mass, we use the $2+1+1$ Lattice QCD FLAG averages FlavourLatticeAveragingGroup:2019iem
$\overline{m}_{s}(2\;{\rm GeV})=(93.44\pm 0.68)\;{\rm MeV}\,.$ (16)
As discussed, the choice of the charm-quark mass definition is more challenging. Since our aim is to get an estimate for the attainable experimental precision only, we do not discuss in detail the different mass schemes for the quarks. For our study, we use the $\overline{\rm MS}$ definition, obtained from $2+1+1$ Lattice QCD FLAG average FlavourLatticeAveragingGroup:2019iem
$\overline{m}_{c}(\overline{m}_{c})=(1.280\pm 0.013)\;{\rm GeV}\ .$ (17)
We treat both masses as external inputs and fixed parameters. We note that this choice gives rather large $\alpha_{s}$ corrections, although the convergence of the series can only be checked by including higher order terms in $\alpha_{s}$. Comparing to the kinetic scheme with $\mu=0.5\,\mathrm{GeV}$ and $m_{c}^{\rm kin}=1.4\,\mathrm{GeV}$, we find that the $\alpha_{s}$ contributions for the lepton energy moments are smaller for the kinetic scheme. Its contribution is further reduced by including also the scheme change of the HQE parameters (as done in Gambino:2010jz ). For the $q^{2}$ moments, we however observe large $\alpha_{s}$ corrections in the kinetic scheme when including the perturbative contribution to $\rho_{D}^{3}$, which should be investigated further. Alternatively, $m_{c}$ and $m_{s}/m_{c}$ could be obtained from the fit (see discussion for $B\to X_{c}$ decays in Gambino:2013rza ).
For the coupling constant, we use
$\alpha_{s}(\overline{m}_{c})=0.386\,,$ (18)
obtained from RunDec with $n_{f}=3$ active flavours Chetyrkin:2000yt . Recently, Wu:2024jyf used the inclusive charm branching ratios together with input for the HQE parameters from the $B\to X_{c}\ell\bar{\nu}_{\ell}$ decays to extract $\alpha_{s}$ employing the kinetic scheme for the charm mass. Our input value is in good agreement with the value extracted in that way. In a similar way as in Wu:2024jyf , a future moments analysis could also allow the determination of the value of the strong coupling constant directly from the data.
The fit is performed in the following way: We define the $\chi^{2}$ function as
$\chi^{2}=\chi^{2}(\mu_{\pi}^{2},\mu_{G}^{2},\rho_{D}^{3},\rho_{LS}^{3},\tau_{0})\textasciiacute\,,$ (19)
depending only on the HQE parameters up to $\mathcal{O}(1/m_{b}^{3})$ and the weak annihilation coefficient $\tau_{0}$. The experimental data used in the fit are the fully inclusive ($q^{2}_{\rm th}=0\,\text{GeV}^{2}$) $\langle q^{2n}\rangle$ and $\langle E_{\ell}^{n}\rangle$ moments with $n=\{1,2,3,4\}$. We take into account the full statistical and systematic covariance between the two sets of spectral moments and the different orders. We perform the fit to the simulated data, construct an Asimov data set Cowan:2010js based on the fit result, and refit the Asimov data to determine the experimental precision. To take into account backgrounds from wrongly reconstructed tags, the statistical uncertainty is scaled by $1/\sqrt{\pi}$, where $\pi$ is the estimated purity based on the previous BESIII inclusive semileptonic analyses CLEO:2009uah ; BESIIIDsIncSL ; BESIII:2022cmg .
The resulting estimates on the attainable absolute uncertainty on the extracted HQE parameters are given in Table 3 and shown in Figure 8. We obtain similar uncertainties when using $m_{c}$ in the kinetic scheme. These estimates show an excellent experimental precision and highlight the feasibility of the suggested program at BESIII. In particular, the experimental uncertainty on the $\rho_{D}^{3}$ extraction looks very promising. The precision on $\mu_{G}^{2}$ may be further improved by putting an external constraint from the mass splitting. As discussed, the uncertainty attainable on the weak annihilation parameter $\tau_{0}$ is challenging to compare with the previous extraction in Gambino:2010jz due the mixing with $\rho_{D}^{3}$. A simple conversion based on the phase-space factors allows an estimate of the improvement. We find over a factor of two up to an order of magnitude improvement. In Gambino:2010jz , the WA parameters were found to be consistent with zero. It will be important to determine if they remain consistent with zero even with improved precision. For the $\Lambda_{c}$ decays, no HQE parameters have been extracted before. For $\rho_{LS}^{3}$, which is only constrained through the lepton energy moments, we note that this parameter has only a small prefactor. Due to RPI, it does not enter the $q^{2}$ moments nor the total rate.
Comparing with the relative uncertainty on the measured moments in Fig. 7, we note that moments of the $D_{s}$ and $\Lambda_{c}$ modes are determined least precise. This is also reflected in the estimated precision on the extracted HQE parameters. We highlight the excellent prospects for the extractions from $D^{+}$ decays due to its large branching fraction. We note that the inclusion of $\alpha_{s}^{2}$ corrections or a different mass definition would shift the central values of the theoretical expressions and might also affect the absolute uncertainty on the extraction. A similar note holds for including a theoretical uncertainty, which allows more flexibility for the fit to accommodate the measured moments. As such, the exact values of the estimated uncertainties should be taken with caution and mainly identify points of improvement for a future experimental analysis.
The uncertainty budget of the $D^{0}$, $D^{+}$, $D_{s}$, and $\Lambda_{c}$ is given in Tables 4, 5, 6, and 7, respectively. In each decay mode, the statistical uncertainty can be improved with a larger BESIII data set, and the calibration uncertainty (labeled as MC stat.) can be improved with increases in the integrated luminosity of the simulated data set. The track reconstruction efficiency is also a leading uncertainty. We stress that the $D_{s}$ and $\Lambda_{c}$ modes specifically would benefit from more precise measurements of the exclusive branching ratios. We show the evolution of the precision estimates in Figure 9, where we assume that all uncertainties will scale with the square root of the integrated luminosities and that BESIII will keep 40 times the integrated luminosity for their simulation with respect to the recorded data. Increased samples of $D_{s}$ mesons and $\Lambda_{c}$ baryons at BESIII would allow for competitive precision on the HQE parameters for all four flavors of charmed hadrons.
Figure 8: Estimated precision on the HQE and weak annihilation parameters in the four decays considered from the simulation described in the text. The error bars include all systematic uncertainties described in the text.
Figure 9: Estimated precision on the HQE and weak annihilation parameters with increased integrated luminosity relative to the current precision for $D^{+}$. We assume that all uncertainties considered scale with $1/\sqrt{\mathcal{L}}$ in the given range.
Table 3: Estimated absolute precision ×$\times 10^{3}$ on the HQE and weak annihilation parameters in the four decays considered from the simulation described in the text. The quoted uncertainty includes all systematic uncertainties considered, and a more detailed separation into individual sources of uncertainties can be found in the text.
$\mu_{\pi}^{2}$ $\mu_{G}^{2}$ $\rho_{D}^{3}$ $\rho_{LS}^{3}$ $\tau_{0}$
$D^{0}$ $21.4$ $29.1$ $1.4$ $37.5$ $66.1$
$D^{+}$ $8.5$ $4.5$ $0.3$ $6.7$ $5.3$
$D_{s}$ $94.5$ $101.1$ $3.1$ $92.9$ $94.4$
$\Lambda_{c}$ $50.8$ $2.8$ $86.1$
Table 4: Absolute uncertainties on the HQE and weak annihilation parameters for the $D^{0}$ channel. A description of the systematic uncertainties can be found in the main text. The values are multiplied by ×$\times 10^{3}$.
$\mu_{\pi}^{2}$ $\mu_{G}^{2}$ $\rho_{D}^{3}$ $\rho_{LS}^{3}$ $\tau_{0}$
Full 21.35 29.13 1.39 37.47 66.12
Stat. 9.17 5.88 0.22 9.63 10.91
MC Stat. 14.69 16.37 0.30 22.64 27.08
$\epsilon_{\mathrm{track.}}$ 16.09 26.81 1.24 32.91 61.95
$\sigma_{\mathrm{track.}}$ 0.22 0.24 0.01 0.33 0.28
$\epsilon_{K_{S}^{0}}$ 1.79 0.77 0.01 1.60 0.94
$\sigma_{K_{S}^{0}}$ 0.04 0.03 0.00 0.03 0.05
$\epsilon_{\gamma}$ 0.72 0.81 0.06 1.13 1.69
$\sigma_{\gamma}$ 0.31 0.08 0.03 0.10 0.87
PID 0.03 0.13 0.00 0.07 0.14
$\mathcal{B}(D^{0}\to K_{1}^{-}\ell\nu_{\ell})$ 5.07 6.02 0.03 8.07 9.60
$\mathcal{B}(D^{0}\to K^{*-}\ell\nu_{\ell})$ 4.06 3.39 0.55 8.72 20.99
$\mathcal{B}(D^{0}\to K^{-}\ell\nu_{\ell})$ 3.64 4.50 0.21 5.38 4.95
$\mathcal{B}(D^{0}\to\rho^{-}\ell\nu_{\ell})$ 6.01 12.61 0.74 13.38 30.13
$\mathcal{B}(D^{0}\to(K\pi)_{S-\mathrm{wave}}\ell\nu_{\ell})$ 6.82 5.63 0.13 9.06 8.89
Table 5: Absolute uncertainties on the HQE and weak annihilation parameters for the $D^{+}$ channel. A description of the systematic uncertainties can be found in the main text. The values are multiplied by ×$\times 10^{3}$.
$\mu_{\pi}^{2}$ $\mu_{G}^{2}$ $\rho_{D}^{3}$ $\rho_{LS}^{3}$ $\tau_{0}$
Full 8.49 4.47 0.30 6.72 5.28
Stat. 4.99 0.85 0.06 3.34 1.59
MC Stat. 3.14 1.42 0.07 2.00 2.71
$\epsilon_{\mathrm{track.}}$ 2.28 2.91 0.20 3.56 1.07
$\sigma_{\mathrm{track.}}$ 0.17 0.03 0.00 0.12 0.06
$\epsilon_{K_{S}^{0}}$ 3.24 1.66 0.15 2.08 1.92
$\sigma_{K_{S}^{0}}$ 0.07 0.06 0.00 0.05 0.04
$\epsilon_{\gamma}$ 1.53 0.93 0.04 0.95 0.94
$\sigma_{\gamma}$ 0.51 0.25 0.04 0.28 1.10
PID 0.03 0.03 0.00 0.04 0.02
$\mathcal{B}(D^{+}\to\eta^{\prime}\ell\nu_{\ell})$ 0.40 0.07 0.02 0.52 1.05
$\mathcal{B}(D^{+}\to\eta\ell\nu_{\ell})$ 1.40 0.32 0.05 1.39 2.17
$\mathcal{B}(D^{+}\to K^{*-}\ell\nu_{\ell})$ 2.63 0.61 0.04 1.68 0.49
$\mathcal{B}(D^{+}\to K^{-}\ell\nu_{\ell})$ 2.18 0.86 0.06 2.24 0.47
$\mathcal{B}(D^{+}\to\omega\ell\nu_{\ell})$ 0.60 0.89 0.08 0.16 0.93
$\mathcal{B}(D^{+}\to\pi^{-}\ell\nu_{\ell})$ 1.60 1.28 0.04 0.52 2.34
$\mathcal{B}(D^{+}\to\rho^{-}\ell\nu_{\ell})$ 0.95 1.05 0.06 1.33 0.65
$\mathcal{B}(D^{+}\to(K\pi)_{S-\mathrm{wave}}\ell\nu_{\ell})$ 4.11 1.75 0.04 1.47 1.32
Table 6: Absolute uncertainties on the HQE and weak annihilation parameters for the $D_{s}$ channel. A description of the systematic uncertainties can be found in the main text. The values are multiplied by ×$\times 10^{3}$.
$\mu_{\pi}^{2}$ $\mu_{G}^{2}$ $\rho_{D}^{3}$ $\rho_{LS}^{3}$ $\tau_{0}$
Full 94.54 101.15 3.05 92.91 94.44
Stat. 34.58 9.63 0.64 28.61 18.03
MC Stat. 25.55 25.78 1.06 21.32 15.39
$\epsilon_{\mathrm{track.}}$ 22.50 28.12 0.96 27.57 21.09
$\sigma_{\mathrm{track.}}$ 8.43 10.08 0.53 5.64 4.20
$\epsilon_{K_{S}^{0}}$ 7.98 6.90 0.22 6.68 12.05
$\sigma_{K_{S}^{0}}$ 8.56 10.05 0.52 5.69 4.22
$\epsilon_{\gamma}$ 7.90 4.28 0.24 11.25 2.37
$\sigma_{\gamma}$ 10.19 7.87 0.56 8.82 7.07
PID 8.56 10.03 0.52 5.68 4.21
$\mathcal{B}(D_{s}\to\eta^{\prime}\ell\nu_{\ell})$ 54.18 9.78 0.21 27.81 10.61
$\mathcal{B}(D_{s}\to\eta\ell\nu_{\ell})$ 33.33 90.26 2.34 60.96 86.46
$\mathcal{B}(D_{s}\to f_{0}\ell\nu_{\ell})$ 14.43 3.31 0.16 8.57 3.12
$\mathcal{B}(D_{s}\to K^{*-}\ell\nu_{\ell})$ 30.21 21.02 0.96 30.77 5.91
$\mathcal{B}(D_{s}\to K^{-}\ell\nu_{\ell})$ 34.83 12.75 0.21 26.67 15.53
$\mathcal{B}(D_{s}\ \to\phi\ell\nu_{\ell})$ 33.44 19.63 0.22 36.21 23.75
Table 7: Absolute uncertainties on the HQE and weak annihilation parameters for the $\Lambda_{c}$ channel. A description of the systematic uncertainties can be found in the main text. The values are multiplied by ×$\times 10^{3}$.
$\mu_{\pi}^{2}$ $\rho_{D}^{3}$ $\tau_{0}$
Full 50.76 2.79 86.11
Stat. 37.04 1.43 46.08
MC Stat. 27.61 0.86 24.45
$\epsilon_{\mathrm{track.}}$ 7.19 0.55 17.38
$\sigma_{\mathrm{track.}}$ 7.41 0.31 10.01
$\epsilon_{K_{S}^{0}}$ 7.21 0.31 9.85
$\sigma_{K_{S}^{0}}$ 6.92 0.31 9.75
$\epsilon_{\gamma}$ 4.42 0.19 6.28
$\sigma_{\gamma}$ 6.04 0.15 4.85
PID 7.21 0.31 9.85
$\mathcal{B}(\Lambda_{c}\to\Delta^{0}\overline{K}^{0}\ell\nu_{\ell})$ 18.63 1.51 46.83
$\mathcal{B}(\Lambda_{c}\to\Lambda^{0}\ell\nu_{\ell})$ 10.29 0.67 17.19
$\mathcal{B}(\Lambda_{c}\to n^{0}\ell\nu_{\ell})$ 15.96 1.55 51.29
$\mathcal{B}(\Lambda_{c}\to\Sigma^{0}\pi^{0}\ell\nu_{\ell})$ 7.65 0.44 17.09
$\mathcal{B}(\Lambda_{c}\to p^{+}K^{-}\ell\nu_{\ell})$ 15.28 1.42 40.80
6 Outlook
As demonstrated, a future program of the study of inclusive charm decays at BESIII is extremely promising. Such a program will improve measurements of the lepton energy moments and allow for the first determinations of the reparametrization invariant $q^{2}$ moments of the kinematical spectrum of $D^{0},D^{+},D_{s}$ and $\Lambda_{c}$ inclusive decays. This opens the way for studying the HQE for the first time in baryonic decays and making a direct comparison between different initial state mesons. Our analysis shows that such a measurement program is achievable, and furthermore highlights how the measurements could further improve. The $D_{s}^{+}$ and $\Lambda_{c}$ measurements have particularly strong prospects for improvement with more data.
We have performed an exploratory analysis of the estimated moments, obtaining a promising attainable precision on the $\{\mu_{\pi}^{2},\mu_{G}^{2},\rho_{D}^{3},\rho_{LS}^{3}\}$ HQE elements and the important weak annihilation $\tau_{0}$ parameter. The precision on the extracted HQE parameters could be further improved by analyzing moments of the spectrum at different kinematic cuts on the $q^{2}$ or lepton energy, in a similar fashion to the $B\to X_{c}\ell\bar{\nu}_{\ell}$ analysis. In addition, moments of the hadronic invariant mass $M_{X}$ could be studied. For this work, we extracted normalized raw moments of the spectrum. In a future analysis, it would also be interesting to compare the theoretical or experimental benefits of normalized raw versus centralized moments.
Combining the extracted HQE elements with measurements of the inclusive branching ratio allows for the extraction of $|V_{cs}|$ and $|V_{cd}|$ for the first time from inclusive decays. The latter requires separating the $c\to s$ from $c\to d$ transition, which as outlined seems feasible in a future analysis. Using the determined HQE parameters, we obtain an experimental precision for $|V_{cs}|$ of 3.3% for $D^{0}$ and $D^{+}$ (with a branching fraction uncertainty contribution of 1.1% CLEO:2009uah ), and 3.8% for $D_{s}^{+}$ (1.3% due to the branching fraction BESIIIDsIncSL ). Given the long-standing puzzle between inclusive and exclusive $|V_{cb}|$, it would be charming to obtain inclusive $|V_{cs}|$ and $|V_{cd}|$ from data.
The HQE parameters have been extracted with good precision from semileptonic $B\to X_{c}\ell\bar{\nu}$ decays, but have never been studied in inclusive charm decays. Extracting the HQE parameters with high precision requires improvements in theory. At minimum, the NNLO and possibly N3LO corrections should be included, and a consistent charm mass definition should be employed. This would then allow for the establishment of a theoretical framework analogous to that used in $B\to X_{c}\ell\bar{\nu}_{\ell}$ decays. In the future, this framework could be merged with the open-source package Kolya which provides theoretical expressions for inclusive $B$ decays Kolya ; kolyapaper .
Comparing the HQE parameters among the different charm species will provide the first test of $SU(3)$ within the HQE and allow valence and non-valence weak annihilation parameters to be separated. Whether or not a HQE analysis of the experimental moments converges will provide an additional and valuable experimental test of the validity of the HQE for charm. Moreover, comparing the HQE parameters between beauty and charm allows for a further test of non-perturbative QCD among different generations. In the infinite mass limit, the HQE parameters in beauty in charm are equal, but in the standard HQE approach they depend in a non-trivial way on the mass of the initial decaying hadron. In Refs. Fael:2019umf ; Gambino:2010jz , the difference between beauty and charm elements is estimated. With the outlined program, these relations can be tested with data.
In addition to these comparisons among different quark species, determining the HQE elements themselves is important as they enter the theoretical predictions for charm lifetimes. Furthermore, the weak annihilation parameters also enter $B\to X_{s,d}\ell\ell$ and $B\to X_{u}\ell\bar{\nu}_{\ell}$ predictions. In this way, the envisioned charm program reaches across different families and would allow for compatibility checks between hadron sectors within the Standard Model. There is a grandeur in this view of semileptonic decays.
Acknowledgements
The work of KKV is supported in part by the Dutch Research Council (NWO) as part of the project Solving Beautiful Puzzles (VI.Vidi.223.083) of the research programme Vidi. The work of FB is supported by DFG Emmy-Noether Grant No. BE 6075/1-1 and BMBF Grant No. 05H21PDKBA The work of MP is supported by the German Research Foundation (DFG) Emmy-Noether Grant No. 526218088. SM and GW are grateful for support from the UK Science and Technology Facilities Council.
Appendix A Calibration
The linear calibration functions for the first, second, third, and fourth $q^{2}$ and $E_{\ell}$ moments of $D^{0}$, $D^{+}$, $D_{s}^{+}$ and $\Lambda_{c}^{2}$ are depicted in Figures. 10-13. The corresponding $C_{\mathrm{calib}}$ and $C_{\mathrm{gen}}$ corrections for all four states factors are depicted in Figures 14 and 15.
Figure 10: $D^{0}$ raw moment calibration
Figure 11: $D^{+}$ raw moment calibration
Figure 12: $D_{s}$ raw moment calibration
Figure 13: $\Lambda_{c}$ raw moment calibration
Figure 14: $\mathcal{C}_{\mathrm{calib}}$ calibration factors.
Figure 15: $\mathcal{C}_{\mathrm{gen}}$ calibration factors.